# Amortized Sampling with Transferable Normalizing Flows

### v2

@@ F.4 JSD evaluation
Model $\downarrow$ $\mathcal{T}\text{‑}\mathrm{JSD}$ $\downarrow$ $\mathcal{T}\text{‑}\mathrm{JSD}$ $\downarrow$ $\mathrm{TICA}\text{‑}\mathrm{JSD}$ $\downarrow$ $\mathcal{T}\text{‑}\mathrm{JSD}$ $\downarrow$ $\mathrm{TICA}\text{‑}\mathrm{JSD}$ $\downarrow$
@@
F.5 Sequence-length extrapolation
YQNPDGSQA (9AA) Chignolin (10AA)
UniSim $10^{5}$ 6.00 0.84 267.68 6.48 0.20
BioEmu 160.52 4.52 1.14 198.90 5.14 0.65
Prose Proposal $10^{9}$ 3.91 1.65 $10^{9}$ 3.63 0.96
Prose SNIS 0.0049 23.79 3.85 1.94 0.0001 832.59 4.35 1.25
Prose (self-improve) Proposal $10^{9}$ 3.73 1.95 $10^{9}$ 3.94 1.13
F.6 Effective sample per second
@@
F.7 Octopeptide Ramachandran plots
@@ F.8 Temperature Plots
Figure 11: TICA plots for RLMM for different temperatures. Reference molecular dynamics (top row), Prose proposal (middle row), Prose SNIS (bottom row) with $5\cdot 10^{4}$ samples.
Figure 12: Energy histogram plots for RLMM for different temperatures.

### v3

@@ F.5 JSD evaluation
Model $\downarrow$ $\mathcal{T}\text{‑}k\text{-}\mathrm{JSD}$ $\downarrow$ $\mathcal{T}\text{‑}k\text{-}\mathrm{JSD}$ $\downarrow$ $\mathrm{TICA}\text{‑}k\text{-}\mathrm{JSD}$ $\downarrow$ $\mathcal{T}\text{‑}k\text{-}\mathrm{JSD}$ $\downarrow$ $\mathrm{TICA}\text{‑}k\text{-}\mathrm{JSD}$ $\downarrow$
@@
bsectionEffective sample per second
@@
F.6 Octopeptide Ramachandran plots
@@ F.7 Temperature Plops
Figure 11: TICA plots for RLMM at different temperatures. Reference molecular dynamics (top row), Prose proposal (middle row), Prose SNIS (bottom row) with $5\cdot 10^{4}$ samples.
Figure 12: Energy histogram plots for RLMM at different temperatures.